# SMReferences for Appendix

### v3

@@ A.6.1 Proof of Theorem 6(i)
where for the last inequality we apply that $|\hat{\tau}-\tau^{*}|\lesssim\sqrt{s_{r}\frac{\log p}{n}}$ and the constraint (16). For the second term, using an analogous analysis as in the proof of Theorem 3, we get that with probability converging to $1$,
@@
For the third term, using the constraint in the calibration program, we have that with probability converging to $1$,
@@
Under the conditions of Theorem 6(i), with probability converging to $1$, we have that $\|\hat{\bm{\mu}}\|_{2}=O(\sqrt{s_{r}\log p})$.
@@
Following an analogous analysis as in the proof of Lemma 8, we have that
@@
is, with probability converging to $1$, a feasible solution to the program used to construct $\hat{\bm{\mu}}$ by choosing $\eta_{r}\asymp\sqrt{\frac{\log p}{n}}$, so that the desired result follows from an analogous analysis as Lemma 2. ∎
A.6.2 Proof of Theorem 6(ii)
@@
Now that all the entries of $\phi^{\prime}(X_{i}^{\top}\gamma^{*})\tilde{X}_{i}$ are bounded and that $e_{i}$’s are i.i.d. bounded random variables which are independent of $\tilde{\mathbf{X}}$, the desired result follows from an application of Hoeffding’s inequality. ∎
@@
Let $\hat{\tau}^{*}_{\textrm{DCal}}$ be a modification of $\hat{\tau}_{\textrm{DCal}}$ but with $\tilde{\pi}$ replaced by $\pi^{*}$. Then we have the decomposition
$\hat{\tau}_{\textrm{DCal}}-\hat{\tau}^{*}_{\textrm{DCal}}=\frac{1}{n}\sum_{i=1}^{n}\frac{(Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}-\hat{\mu}_{i})(\pi_{i}^{*}-\tilde{\pi}_{i})}{\hat{\sigma}_{e}^{2}}+\left(\frac{1}{\hat{\sigma}_{e}^{2}}-\frac{1}{\bar{\sigma}_{e}^{2}}\right)\cdot\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}-\hat{\mu}_{i})(T_{i}-\pi_{i}^{*}).$ (21)
@@
For the first term, using Taylor expansion followed by a remainder control (see e.g. the analysis of term $\mathrm{I}$ in the proof of Theorem 5), it remains to show that
$\mathrm{I}\coloneqq\left|\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}-\hat{\mu}_{i})\phi^{\prime}(X_{i}^{\top}\gamma^{*})\tilde{X}_{i}^{\top}(\tilde{\gamma}-\gamma^{*})\right|=o_{\mathbb{P}}(1/\sqrt{n}).$
To achieve this goal, using mean value theorem that $\phi^{\prime}(X_{i}^{\top}\gamma^{*})-\phi^{\prime}(X_{i}^{\top}\hat{\gamma})=\phi^{\prime\prime}(\eta_{i}X_{i}^{\top}\gamma^{*}+(1-\eta_{i})X_{i}^{\top}\hat{\gamma})$ for some $\eta_{i}\in[0,1]$, and that $\phi^{\prime\prime}(\cdot)$ is a uniformly bounded function, we have that
$\mathrm{I}\lesssim\left|\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}-\hat{\mu}_{i})\phi^{\prime}(X_{i}^{\top}\hat{\gamma})\tilde{X}_{i}^{\top}(\tilde{\gamma}-\gamma^{*})\right|+\frac{1}{n}\sum_{i=1}^{n}\left|(\tilde{X}_{i}^{\top}(\tilde{\gamma}-\gamma^{*}))(X_{i}^{\top}(\hat{\gamma}-\gamma^{*}))\right|.$
Now applying Hölder’s inequality to the first term and Cauchy-Schwarz inequality to the second, we further have
$\mathrm{I}\lesssim$ $\left\|\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}-\hat{\mu}_{i})\phi^{\prime}(X_{i}^{\top}\hat{\gamma})\tilde{X}_{i}\right\|_{\infty}\|\tilde{\gamma}-\gamma^{*}\|_{1}+\frac{1}{n}\left\|\tilde{\mathbf{X}}(\tilde{\gamma}-\gamma^{*})\right\|_{2}\left\|{\mathbf{X}}(\hat{\gamma}-\gamma^{*})\right\|_{2}.$
With Lemma 13, following analysis analogous to the third term in the decomposition (19), we have that the second term in the above inequality is of order $o_{\mathbb{P}}(1/\sqrt{n})$. For the first term, using again Lemma 13 and Condition 8, we have that with probability converging to $1$, $\|\tilde{\gamma}-\gamma^{*}\|_{1}\lesssim s_{\pi}\sqrt{\frac{\log p}{n}}$. Thus, to prove that term $\mathrm{I}$ is of order $o_{\mathbb{P}}(1/\sqrt{n})$, it remains to show that with probability converging to $1$,
$\left\|\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}-\hat{\mu}_{i})\phi^{\prime}(X_{i}^{\top}\hat{\gamma})\tilde{X}_{i}\right\|_{\infty}\lesssim\sqrt{\frac{\log p\vee n}{n}}.$
The above result then follows exactly the same lines of proof as in (20).
@@
$\sqrt{n}(\hat{\tau}_{\textrm{DCal}}-\tau^{\ast})=\sqrt{n}(\hat{\tau}_{\textrm{DCal}}^{*}-\tau^{\ast})+o_{\mathbb{P}}(1).$
@@
$\sqrt{n}(\hat{\tau}^{*}_{\textrm{DCal}}-\tau^{*})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{e_{i}\varepsilon_{i}}{\sigma_{e}^{2}}+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{e_{i}((\tau^{*}-\hat{\tau})\pi_{i}^{*}+r_{i}^{*}-\hat{r}_{i}-\hat{\mu}_{i})}{\sigma_{e}^{2}}+o_{\mathbb{P}}(1)$
@@
Reorganizing $Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}=r_{i}^{*}+e_{i}(\tau^{*}-\hat{\tau})+\pi_{i}^{*}(\tau^{*}-\hat{\tau})+r_{i}^{*}-\hat{r}_{i}+\varepsilon_{i}$, we rewrite $\hat{\tau}^{*}_{\textrm{DCal}}$ as
$\hat{\tau}^{*}_{\textrm{DCal}}=\tau^{*}+\frac{1}{n}\sum_{i=1}^{n}\frac{e_{i}((\tau^{*}-\hat{\tau})\pi_{i}^{*}+r_{i}^{*}-\hat{r}_{i}-\hat{\mu}_{i}+\varepsilon_{i})}{\bar{\sigma}_{e}^{2}}.$
@@
$\frac{1}{n}\sum_{i=1}^{n}e_{i}((\tau^{*}-\hat{\tau})\pi_{i}^{*}+r_{i}^{*}-\hat{r}_{i}-\hat{\mu}_{i})=O_{\mathbb{P}}\left(\frac{1}{\sqrt{n}}\right).$
@@
Appendix B Proof of Corollary 7
Similar to Section 2 and 3, the proof of Corollary 7 can be divided into the following steps. First, under $\xi_{\pi}>1/2$, $\hat{\tau}_{{\textrm{SCal},r}}$ is $\sqrt{n}$-consistent for estimating ATE.
@@ Appendix B Proof of Corollary 7
Under the conditions of Corollary 7(i), we have the following:
$\sqrt{n}(\hat{\tau}_{{\textrm{SCal},r}}-\bar{\tau}^{\ast})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{T_{i}\varepsilon_{i}(1)}{\pi_{i}^{\ast}}+o_{\mathbb{P}}(1).$
@@
The proof essentially follows from the proof of Theorem 2 of Wang and Shah [2020], by using the constraint (5) and Condition 13. Specifically, we bound the estimation error as follows:
$\hat{\tau}_{{\textrm{SCal},r}}-\bar{\tau}^{\ast}$ $=\frac{1}{n}\sum_{i=1}^{n}\frac{T_{i}}{\hat{\pi}_{i}}(Y_{i}-\hat{r}_{i}-\hat{\mu}_{i})-\frac{T_{i}}{\phi(X_{i}^{\top}\gamma^{\ast})}(Y_{i}-r_{i}^{\ast})+\hat{r}_{i}+\hat{\mu}_{i}-r_{i}^{\ast}$
$=\frac{1}{n}\sum_{i=1}^{n}T_{i}\left(\frac{1}{\phi(X_{i}^{\top}\hat{\gamma})}-\frac{1}{\phi(X_{i}^{\top}\gamma^{\ast})}\right)(Y_{i}-\hat{r}_{i}-\hat{\mu}_{i})+\left(1-\frac{T_{i}}{\phi(X_{i}^{\top}\gamma^{\ast})}\right)(\hat{r}_{i}+\hat{\mu}_{i}-r_{i}^{\ast})$
where the first term is upper bounded by $o(n^{-1/2})$ using constraint (5) and Condition 13 and the second term is also negligible because $\phi(x^{\top}\gamma^{\ast})$ approximates $\pi^{\ast}(x)$ at rate much smaller than $n^{-1}$ in $L_{2}(\mathbb{P}_{\theta^{\ast}})$ norm by taking $p\gtrsim n^{2}$. ∎
Next, we follow the analysis strategy of Section 3.1. As a first step, we need to show the following.
@@
The same conclusion of Lemma 2 holds when Condition 7 is replaced by Condition 12 and $s_{r}=o(\sqrt{n}/\log p)$ is replaced by $\xi_{r}>1/2$.

### v4

@@ B.1.1 Proof of Theorem 6 (i)
where for the last inequality we apply that $|\hat{\tau}-\tau^{*}|\lesssim\sqrt{s_{r}\frac{\log p}{n}}$ and the constraint (16). For the second term, following again the constraint (16), we get that with probability converging to $1$,
$\mathrm{II}\lesssim\frac{1}{n}\sum_{i=1}^{n}(T_{i}-\tilde{\pi}(X_{i}))(r(X_{i})-\hat{r}(X_{i})).$
From above and applying exactly the same proof as Theorem 3, with probability converging to $1$,
@@
For the third term, it directly follows from the constraints in the calibration program that with probability converging to $1$,
@@
Consider the set up of Theorem 6 (i), with probability converging to $1$, we have that $\|\hat{\bm{\mu}}\|_{2}=O(n^{1/4})$.
@@
Following the same analysis as Lemma 8, we have that by choosing $\eta_{r}\asymp\sqrt{\frac{\log p}{n}},M_{r}\asymp 1$ large enough,
@@
is, with probability converging to $1$, a feasible solution to the optimization program, so that the desired result follows from an analysis analogous to Lemma 2. ∎
B.1.2 Proof of Theorem 6 (ii)
@@
Now that all the entries of $\phi^{\prime}(X_{i}^{\top}\gamma^{*})\tilde{X}_{i}$ are bounded and that $e_{i}$’s are i.i.d. bounded random variables which are independent from $\tilde{\mathbf{X}}$, the desired result follows from an application of Hoeffding’s inequality. ∎
@@
Let $\hat{\tau}^{*}_{\mathrm{DCal}}$ be a modification of $\hat{\tau}_{\mathrm{DCal}}$ but with $\tilde{\pi}$ replaced by $\pi^{*}$. Then we have the decomposition
$\hat{\tau}_{\mathrm{DCal}}-\hat{\tau}^{*}_{\mathrm{DCal}}=\frac{1}{n}\sum_{i=1}^{n}\frac{(Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}-\hat{\mu}_{i})(\pi_{i}^{*}-\tilde{\pi}_{i})}{\hat{\sigma}_{e}^{2}}+\left(\frac{1}{\hat{\sigma}_{e}^{2}}-\frac{1}{\bar{\sigma}_{e}^{2}}\right)\cdot\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}-\hat{\mu}_{i})(T_{i}-\pi_{i}^{*}).$ (31)
@@
For the first term, recall we already have the constraint (16), it remains to show that
$\mathrm{I}\coloneqq\left|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}-\hat{\mu}_{i})(\pi^{*}_{i}-\tilde{\pi}_{i})\right|=o_{\mathbb{P}}(1).$
With Lemma 12, this directly follows from the exactly the same analysis as the terms $\mathrm{I}$ and $\mathrm{II}$ in the proof of Theorem 5.
@@
$\sqrt{n}(\hat{\tau}_{\mathrm{DCal}}-\tau^{\ast})=\sqrt{n}(\hat{\tau}_{\mathrm{DCal}}^{*}-\tau^{\ast})+o_{\mathbb{P}}(1).$
@@
$\sqrt{n}(\hat{\tau}^{*}_{\mathrm{DCal}}-\tau^{*})=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{e_{i}\varepsilon_{i}}{\sigma_{e}^{2}}+\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\frac{e_{i}((\tau^{*}-\hat{\tau})\pi_{i}^{*}+r_{i}^{*}-\hat{r}_{i}-\hat{\mu}_{i})}{\sigma_{e}^{2}}+o_{\mathbb{P}}(1)$
@@
Reorganizing $Y_{i}-T_{i}\hat{\tau}-\hat{r}_{i}=e_{i}(\tau^{*}-\hat{\tau})+\pi_{i}^{*}(\tau^{*}-\hat{\tau})+r_{i}^{*}-\hat{r}_{i}+\varepsilon_{i}$, we rewrite $\hat{\tau}^{*}_{\mathrm{DCal}}$ as
$\hat{\tau}^{*}_{\mathrm{DCal}}=\tau^{*}+\frac{1}{n}\sum_{i=1}^{n}\frac{e_{i}((\tau^{*}-\hat{\tau})\pi_{i}^{*}+r_{i}^{*}-\hat{r}_{i}-\hat{\mu}_{i}+\varepsilon_{i})}{\bar{\sigma}_{e}^{2}}.$
@@
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}e_{i}((\tau^{*}-\hat{\tau})\pi_{i}^{*}+r_{i}^{*}-\hat{r}_{i}-\hat{\mu}_{i})=O_{\mathbb{P}}(1).$
@@
B.2 Conditions related to Section 4.2
In this section, we lay out the standard identification conditions for $\chi^{\ast}$ defined in (17) to identify LATE, i.e. $\chi^{\ast}\equiv\mathbb{E}[Y(1)-Y(0)|T(1)>T(0)]$. This is now a classical result in the causal inference literature, and hence we will not prove it or discuss it further.
Condition 13 (IV identification conditions).
  1. (i)
[…]
  2. (ii)
    Exclusion restriction: $Y(1,t)=Y(0,t)=Y(t)$ almost surely, for all $t\in\{0,1\}$;
  3. (iii)
    Monotonicity: $T(1)\geq T(0)$ almost surely;
  4. (iv)
    Relevance: $\mathbb{P}(T(1)=1)\neq\mathbb{P}(T(0)=1)$;
  5. (v)
    Consistency: $Y=TY(1)+(1-T)Y(0)$ and $T=ZT(1)+(1-Z)T(0)$;
  6. (vi)
    IV Positivity: There exists an absolute constant $c_{\zeta}\in(0,0.5)$ such that $\zeta^{\ast}(X)\in(c_{\zeta},1-c_{\zeta})$ almost surely.
B.3 Proof of Corollary 7
Similar to Section 2 and 3, the proof of Corollary 7 can be divided into the following steps as in the proof of Theorem 3 and 5.
As a first step, we need to show the following.
@@ B.3 Proof of Corollary 7
The same conclusion of Lemma 2 holds when Condition 6 is replaced by Condition 11 and $s_{r}=o(\sqrt{n}/\log p)$ is replaced by $\xi_{r}>1/2$.
@@
The proof follows a similar argument to the proof of Lemma 2 in Appendix A.2, except that (22) needs to be replaced by the following chain of inequalities:
$\|\hat{\bm{\mu}}\|_{2}$ $\leq\|\hat{\bm{\mu}}_{\texttt{ora}}\|_{2}\lesssim\sqrt{\sum_{i=1}^{n}(\hat{r}_{i}-r_{i}^{\ast})^{2}}\lesssim\sqrt{\sum_{i=1}^{n}(\bar{r}_{i}^{\ast}-r_{i}^{\ast})^{2}}+\sqrt{\sum_{i=1}^{n}(X_{i}^{\top}(\hat{\beta}-\beta^{\ast}))^{2}}.$
Then the rest follows from applying the same argument as in Appendix A.2, together with the assumption (18) in Condition 10. ∎
We now prove Corollary 7 (i).
Proof.
The proof closely mirrors the proof for Theorem 3 given in Appendix A.3, except that we need to modify the decomposition in (24) with an extra term $\mathrm{IV}$:
$\hat{\tau}_{\mathrm{DCal}}$ $-\bar{\tau}^{\ast}$
$=$ $\ \underset{\mathrm{I}}{\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(1-\frac{T_{i}}{\tilde{\pi}_{i}}\right)\psi^{\prime}(X_{i}^{\top}\hat{\beta})X_{i}^{\top}(\hat{\beta}-\beta^{*})}}$
$+\underset{\mathrm{II}}{\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(1-\frac{T_{i}}{\tilde{\pi}_{i}}\right)\psi^{\prime\prime}(\eta_{i}X_{i}^{\top}\hat{\beta}+(1-\eta_{i})X_{i}^{\top}\beta^{*})\left(X_{i}^{\top}(\hat{\beta}-\beta^{*})\right)^{2}}}$
$+\underset{\mathrm{III}}{\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(1-\frac{T_{i}}{\tilde{\pi}_{i}}\right)\hat{\mu}_{i}}}+\underbrace{\frac{1}{n}\sum_{i=1}^{n}\left(1-\frac{T_{i}}{\tilde{\pi}_{i}}\right)\left(\bar{r}^{\ast}_{i}-r^{\ast}_{i}\right)}_{\mathrm{IV}}+\frac{1}{n}\sum_{i=1}^{n}\frac{T_{i}\varepsilon_{i}(1)}{\tilde{\pi}_{i}}.$
Using the same argument as in Appendix A.3, with Condition 6 replaced by Condition 11, we have, since $\xi_{r}>1/2$,
$|\mathrm{I}|=O_{\mathbb{P}}\left(\left(\sqrt{\frac{n}{\log p}}\right)^{-\frac{4\xi_{r}}{2\xi_{r}+1}}\right)=o_{\mathbb{P}}(n^{-1/2}),|\mathrm{II}|\lesssim\|\hat{\beta}-\beta^{\ast}\|^{2}=O_{\mathbb{P}}\left(\left(\sqrt{\frac{n}{\log p}}\right)^{-\frac{4\xi_{r}}{2\xi_{r}+1}}\right)=o_{\mathbb{P}}(n^{-1/2}).$
Following Lemma 15, we also have $|\mathrm{III}|=o_{\mathbb{P}}(n^{-1/2})$. Finally, we can easily show that $|\mathrm{IV}|=o_{\mathbb{P}}(n^{-1/2})$ by invoking (18) using Cauchy-Schwarz inequality. ∎
Finally, we are left to prove Corollary 7 (ii). To this end, we first need to show that $\gamma^{\ast}$ is a feasible solution to (9a), or equivalently $\bar{\pi}^{\ast}_{i}$’s satisfy all the constraints in (9a), with probability converging to 1.
Proof.
First, we notice as in the proof of Lemma 1, the true propensity score $\pi^{\ast}_{i}$’s satisfy all the constraints in (9a) with probability converging to 1. Hence, after applying triangle inequality and Cauchy-Schwarz inequality, we are only left to show the following holds with probability approaching 1:
$\left\{\frac{1}{n}\sum_{i=1}^{n}\left(\frac{T_{i}}{\pi^{\ast}_{i}}-\frac{T_{i}}{\bar{\pi}^{\ast}_{i}}\right)^{2}\right\}^{1/2}\ll\sqrt{\frac{\log p}{n}},$
which is true due to (19) imposed in Condition 12. ∎
Next, we establish the following lemma analogous to Lemma 4.
@@
Under the Conditions of Lemma 4, except with Condition 7 replaced by Condition 12 and $s_{\pi}=o(\sqrt{n}/\log p)$ replaced by $\xi_{\pi}>1/2$, by choosing $\eta_{r},\eta_{\pi_{1}},\eta_{\pi_{2}}\asymp\sqrt{\frac{\log p}{n}}$, $M_{r},M_{\pi},M_{\gamma}\asymp 1$, we have that with probability converging to 1,
$\|\tilde{\gamma}-\hat{\gamma}\|_{1}=O\left(\left(\sqrt{\frac{n}{\log p}}\right)^{-\frac{2\xi_{\pi}-1}{2\xi_{\pi}+1}}\right)\,\,\,\&\,\,\,\frac{1}{n}\sum_{i=1}^{n}(X_{i}^{\top}(\tilde{\gamma}-\hat{\gamma}))^{2}=O\left(\left(\sqrt{\frac{n}{\log p}}\right)^{-\frac{4\xi_{\pi}}{2\xi_{\pi}+1}}\right).$